# Phase Space Electronic Structure Theory: From Diatomic Lambda-Doubling to Macroscopic Einstein-de Haas

### v1

@@ 5 Ab Initio Results
Let us now address the necessary calculations. Below, we fixed the bond length of NO to be the experimental value NIST_CCCBDB_10102439 1.154 Å and ignored dependence on vibrational and centrifugal distortion. Except where otherwise noted (see Sec. 5.2), we treat SOC with the one-electron Breit-Pauli Hamiltonian bethe2013quantum (not including the two-electron spin-same-orbit and spin-other-orbit coupling). In our study of NO, as a sanity check, the first thing we calculated was the energy splitting between the $^{2}\Pi_{1/2}$ and $^{2}\Pi_{3/2}$ manifold at $L=0$ due to SOC. For these calculations, we have run full configuration interaction (FCI) calculations for absolute certainty. The low-energy eigenstates of NO, illustrated in Fig. 2, are split by strong SOC into the ground state $^{2}\Pi_{1/2}$ manifold and a $^{2}\Pi_{3/2}$ manifold that lies 143 cm-1 higher, in reasonable agreement with the experimental separation james1964spin of 123 cm-1 with about 16% overestimation. Note that this splitting is far bigger than the $\Lambda$-splitting; see Fig. 2.
Next, let us address the $\Lambda$-splitting. As mentioned above, if we wish to use the approach in Sec. 4.3, there are two ways to extract the key objects, $\mathbf{L}^{e}$ and $\mathbf{S}^{e}$ (or really just $\mathbf{L}^{e}$ in practice), from a phase space calculation (and see Table 1 below.).
@@
5.1 Estimate of Coupling From PS calculations
@@
    We solved the same Hamiltonian using the density matrix renormalization group (DMRG) solver zhai2023block2 with enough bond dimension to converge to the FCI accuracy in a slightly larger basis 6-31G, after benchmarking against FCI in the STO-3G basis. For these calculations, the ground and first three excited states were solved using a state-averaged DMRG approach dorando2007targeted. Bond dimensions of 500 and 3000 were used for the STO-3G and 6-31G basis sets, respectively, to converge the first excitation energy to about $10^{-12}$ and $10^{-9}$ Ha precision. Unlike the calculations for (i) and (iii), for better numerical accuracy when evaluating a small energy gap, we invoked a linear approximation for the phase space energies and ran calculations with $L=3\hbar$ instead of $\hbar/2$ and rescaled the final energy splitting by 1/6; the assumption that the $\Lambda$-splitting scales approximately linearly with $J$ for small $J$ follows from the parabolic nature of Fig. 1.
@@
    In order to reach a full basis limit, we developed and implemented a constrained Hartree-Fock (cHF) ma2015constrained, kaduk2012constrained, peng2025accurate, behler2005dissociation approach which avoids the incorrect broken symmetry solutions with $S^{e}_{c}=\pm 1/2$ and $L^{e}_{c}=\mp 1$ as found by standard HF theory. Here, our constraint penalizes finite $\mathbf{S}^{e}$ vectors by adding an energy penalty term $\Delta E=-\lambda|\mathbf{S}^{e}|^{2}$ and therefore yields the correct parity eigenstates with $S^{e}_{c},L^{e}_{c}\approx 0$. We applied this cHF method first with $\lambda=0.1$, and then relaxed the converged solution with $\lambda=0$ until convergence again. To obtain the first excited state with the opposite parity, we used the time reversal of the ground state electronic wavefunction as the initial guess, converged the cHF calculation with $\lambda=0.1$, and then relaxed the solution with $\lambda=0$.
@@ 5.2 The Splitting as a function of $J$
Figure 4: Energy splitting between the $e$ and $f$ states as a function of total angular momentum $J$ from the 2D model parametrized with $\alpha$ derived from FCI in the STO-3G basis (blue) and from cHF in the aug-ccpVTZ basis (red), both using Method III, in comparison to experiments (black). Dashed lines include only the 1-electron SOC, while solid lines include both 1-electron and 2-electron SOC (with a spin-orbit mean-field approximation neese2005efficient, lee2023ab).
@@ 6.1 Kramers’ Theorem
As a result, if one applies the time reversal operator to the total (nuclear + electronic wavefunction in Eq.  4.1), one obtains:
@@
6.2 A Phase Space Visualization of the e/f States
@@
Figure 5: a) Our original view (see Fig. 1) of potential energy surfaces of the ground and first excited states of NO according to a phase space electronic structure Hamiltonian as a function of the canonical nuclear angular momentum. b) The same data now labeled differently: The green and red curves can be understood as the $e$ and $f$ parity total electronic and nuclear states, now understood to functions of total angular momentum. The $e/f$ states are the states that are observed spectroscopically.
The phase space framework above paints a very intuitive picture for visualizing the $e$ and $f$ parity (total electronic and nuclear) wavefunctions in Eq. 4.1. To establish such a picture, let us explore the phase space PES for a rotation around the $a$ axis, e.g. $L_{a}>0$. We rewrite the dominant components of the low-energy eigenstate wavefunctions Eq. 4.1 as
@@
$+\frac{1}{\sqrt{2}}(|J\Omega M\rangle-|J-\Omega M\rangle)\frac{1}{\sqrt{2}}(|\Lambda S-\Sigma\rangle\mp|-\Lambda S\Sigma\rangle)].$
@@
are the two electronic eigenstates that are symmetric with respect to the reflection through the plane $bc$, up to a phase factor. According to Eq. 61, when the molecule has a positive angular momentum around the a-axis, the electronic ground state is dominantly $|\zeta^{+}\rangle$; when the molecule has a negative angular momentum around the a-axis, the electronic ground state is dominantly $|\zeta^{-}\rangle$. This fact implies that, whereas the crossing phase space PESs in Fig. 1 (copied in Fig. 5a) are defined (and labeled) by their constant electronic character, the observed ee/$f$ states arise and are defined (and labeled) by their relative energies as in Fig. 5b. In the latter case, the ee/$f$ states are not so much discrete states but rather arise as part of a continuous potential energy surface that must be requantized (as in Sec. 4.3 above).
@@
From a quantum chemistry perspective, if one considers the two BO electronic states of NO in one orientation as the degenerate "diabatic states" of interest, then the nuclear motion breaks this degeneracy, leading to the "adiabatic states" of interest in Fig. 5. However, the ee/$f$ states have identical energies at L = 0 because of the form of the splitting in Eq. 51, so that there are two different labeling schemes – Fig. 5a versus Fig. 5b. In a sense, one can argue that NO has a “conical intersection” yarkony1996diabolical, baer2006beyond, lehman2014dynamical in momentum space, which provides a new interpretation of $\Lambda$-doubling (and a different view of the magnetic monopole derivative couplings discussed in Refs. moody1986realizations, bohm2013geometric). Note that for the case of a diatomic molecule (as in Eq. 49), the conical intersection in phase space has co-dimension two (because only $L_{a}$ and $L_{b}$ enter the rotational Hamiltonian); more generally, in line with Ref. duston2025conical, we would expect the conical intersection around ${\bf L}=0$ to have co-dimension three. Note also that our analysis above of Eq. 51 is effectively quantitative insofar as a purely quantum calculation using $L_{a}=L_{b}=\frac{1}{2}(J+\frac{1}{2})$ will necessarily lead to an energy splitting proportional to a factor $2\cdot\frac{1}{2}(J+\frac{1}{2})\approx\sqrt{J(J+1)}$, which is in agreement with Eq. 51 above.
@@ 7 Conclusions and Outlook
We have demonstrated that a simple phase space perspective on electronic structure theory provides a natural framework to capture the $\Lambda$-doubling in diatomic molecules. Thus, in line with previous discussionsbian2025review, bradbury2025symmetry, we assert that the same beyond-BO physics that underlies one of the smallest microscopic diatomic phenomena is also responsible for the macroscopic Einstein de-Haas effecteinstein1915experimenteller, mentink2019quantum, ganzhorn2016quantum and, perhaps, chiral-induced spin selectivitybloom2024chiral, ray1999asymmetric, gohler2011spin. Interestingly, the major result here has been the nonadiabatic splitting $\Delta E$ between different states of different angular momentum, where dynamically $\hbar/\Delta E$ corresponds to the tunneling time between degenerate states. In the context of the Einstein de Haas effect, one is less interested in parity eigenstates (e.g. ee/$f$), but rather in a basis with maximally aligned and opposite spin directions along the magnetic field direction. Nevertheless, the approach above should be able to isolate the tunneling time for a spin (or a collection of spins) to change orientation. In other words, whereas most treatments of Einstein de Haas phenomena use angular momentum conservation to predict the steady state nuclear rotational angular momentum that arises upon a total change in electronic spin angular momentum, the present approach based on phase space electronic structure theory will be able to go further and calculate transient effects: How long does it take for the spins to align? For the metal to reach a steady state of rotation? And of course, in the long run, how much energy is lost if we allow for frictional losses from nuclear vibrations?
With regards to this last point, although in this work we have restricted our attention to electron coupling to purely rotational nuclear motion, the most interesting physics undoubtedly lies ahead as we consider larger molecules and materials, and will need to model both vibrations and rotations – and their coupled effect on electrons burroughs2024infrared, davis2023infrared, davis2024bimolecular – all on an equal footing. Interestingly, and in retrospect, from this perspective, the current article is in a sense trivial. After all, our major finding is that, in order to recover $\Lambda$-doubling, all that is required is that (i) one include SOC within a BO calculation and then (ii) one couple together the nuclear and electronic angular momentum through the $\hat{{\bf L}}^{e}$ operator, as in Eq. 42 above. In this sense, in fact, one can recover Fig. 4 without really ever invoking the phase space electronic structure Hamiltonian in Sec. 2; if one wishes to avoid exploring phase space potential energy surfaces, one can model $\Lambda$-doubling directly with BO states and without building the $\hat{\mathbf{\Gamma}}$ operator.
@@
That being said, in the future, if we seek to model larger molecules or materials – any form of matter with more than two degrees of freedom – it will be essential to build the $\hat{\mathbf{\Gamma}}$ operators and decompose the electronic momentum and angular momentum across different nuclear degrees of freedom. Indeed, our strong feeling is that, given the success of the current manuscript in tandem with previous results in VCDduston2024phase and ROAtao2025non, we are now equipped to explore a very broad set of questions: what are the roles of chiral phonons in determining spin-dependent electron transfer rates kim2023chiral? How should we best quantitatively describe spin relaxationlunghi2019phonons? Can we definitively prove the nature of the fundamental mechanisms behind different magnetic field effectssteiner1989magnetic? Lastly, this manuscript has suggested that whenever the ground state has a degeneracy built on different electronic angular momenta, nuclear motion will lead to a small splitting. As such, might not the equations above (especially Eq. 49, which describes a gap opening due to off-diagonal electron-phonon coupling and is analogous to a Su–Schrieffer–Heeger (SSH)/Peierls Hamiltonian in the case of electron hoppingsu1979solitons, su1980soliton) also apply to a metal and be a starting point for superconductivity calculations? When we move beyond a simple diatomic molecule, we can expect the number and the nature of interesting problems involving angular-momentum transfer to grow very rapidly.

### v2

@@ 5 Ab Initio Results
Let us now address the necessary calculations. Below, we fixed the bond length of NO to be the experimental value 31 1.154 Å and ignored dependence on vibrational and centrifugal distortion. Except where otherwise noted (see below), we treat SOC with the one-electron Breit-Pauli Hamiltonian 5 (not including the two-electron spin-same-orbit and spin-other-orbit coupling). In our study of NO, as a sanity check, the first thing we calculated was the energy splitting between the $^{2}\Pi_{1/2}$ and $^{2}\Pi_{3/2}$ manifold at $L=0$ due to SOC. For these calculations, we have run full configuration interaction (FCI) calculations for absolute certainty. The low-energy eigenstates of NO, illustrated in Fig. 2, are split by strong SOC into the ground state $^{2}\Pi_{1/2}$ manifold and a $^{2}\Pi_{3/2}$ manifold that lies 143 cm-1 higher, in reasonable agreement with the experimental separation 30 of 123 cm-1 with about 16% overestimation. Note that this splitting is far bigger than the $\Lambda$-splitting; see Fig. 2.
Next, let us address the $\Lambda$-splitting. In line with the discussion above, we will focus on the splitting of the ground doublet in the $^{2}\Pi_{1/2}$ manifold. As mentioned above, if we wish to use the approach in Sec. 4, there are two ways to extract the key objects, $\mathbf{L}^{e}$ and $\mathbf{S}^{e}$ (or really just $\mathbf{L}^{e}$ in practice), from a phase space calculation (and see Table 1 below.).
@@
Estimate of Coupling From PS calculations
@@
    We solved the same Hamiltonian using the density matrix renormalization group (DMRG) solver 72 with enough bond dimension to converge to the FCI accuracy in a slightly larger basis 6-31G, after benchmarking against FCI in the STO-3G basis. For these calculations, the ground and first three excited states were solved using a state-averaged DMRG approach 18. Bond dimensions of 500 and 3000 were used for the STO-3G and 6-31G basis sets, respectively, to converge the first excitation energy to about $10^{-12}$ and $10^{-9}$ Ha precision. Unlike the calculations for (i) and (iii), for better numerical accuracy when evaluating a small energy gap, we invoked a linear approximation for the phase space energies and ran calculations with $L=3\hbar$ instead of $\hbar/2$ and rescaled the final energy splitting by 1/6; the assumption that the $\Lambda$-splitting scales approximately linearly with $J$ for small $J$ follows from the parabolic nature of Fig. 1.
@@
    In order to reach a full basis limit, we developed and implemented a constrained Hartree-Fock (cHF) 42, 32, 48, 4 approach which avoids the incorrect broken symmetry solutions with $S^{e}_{c}=\pm 1/2$ and $L^{e}_{c}=\mp 1$ as found by standard HF theory. Here, our constraint penalizes finite $\mathbf{S}^{e}$ vectors by adding an energy penalty term $\Delta E=-\lambda|\mathbf{S}^{e}|^{2}$ and therefore yields the correct parity eigenstates with $S^{e}_{c},L^{e}_{c}\approx 0$. We applied this cHF method first with $\lambda=0.1$, and then relaxed the converged solution with $\lambda=0$ until convergence again. To obtain the first excited state with the opposite parity, we used the time reversal of the ground state electronic wavefunction as the initial guess, converged the cHF calculation with $\lambda=0.1$, and then relaxed the solution with $\lambda=0$.
@@
Figure 4: Energy splitting between the $e$ and $f$ states as a function of total angular momentum $J$ from the 2D model parametrized with $\alpha$ derived from FCI in the STO-3G basis (blue) and from cHF in the aug-ccpVTZ basis (red), both using Method III, in comparison to experiments (black). Dashed lines include only the 1-electron SOC, while solid lines include both 1-electron and 2-electron SOC (with a spin-orbit mean-field approximation 46, 36).
@@ 6 Discussion
As a result, if one applies the time reversal operator to the total (nuclear + electronic wavefunction in Eq.  4), one obtains:
@@
A Phase Space Visualization of the e/f States
@@
Figure 5: a) Our original view (see Fig. 1) of potential energy surfaces of the ground and first excited states of NO according to a phase space electronic structure Hamiltonian as a function of the phase space canonical nuclear angular momentum $L$. b) The same data now labeled differently: The green and red curves can be understood as the $e$ and $f$ parity total electronic and nuclear states, now understood as functions of total angular momentum $J$ (represented by the phase space $L$). The $e/f$ states are the states that are observed spectroscopically.
The phase space framework above paints a very intuitive picture for visualizing the $e$ and $f$ parity (total electronic and nuclear) wavefunctions in Eq. 4. To establish such a picture, let us explore the phase space PES for a rotation around the $a$ axis, e.g. $L_{a}>0$. We rewrite the dominant components of the low-energy eigenstate wavefunctions Eq. 4 as
@@
$+\frac{1}{\sqrt{2}}(|J\Omega M\rangle-|J-\Omega M\rangle)\frac{1}{\sqrt{2}}(|\Lambda S-\Sigma\rangle\mp|-\Lambda S\Sigma\rangle)]$
@@
are the two electronic eigenstates that are symmetric with respect to the reflection through the plane $bc$, up to a phase factor. According to Eq. 61, when the molecule has a positive angular momentum around the a-axis, the electronic ground state is dominantly $|\zeta^{+}\rangle$; when the molecule has a negative angular momentum around the a-axis, the electronic ground state is dominantly $|\zeta^{-}\rangle$. This fact implies that, whereas the crossing phase space PESs in Fig. 1 (copied in Fig. 5a) are defined (and labeled) by their constant electronic character, the observed ee/$f$ states arise and are defined (and labeled) by their relative energies as in Fig. 5b. In the latter case, the ee/$f$ states are not so much discrete states but rather arise as part of a continuous potential energy surface that must be requantized (as in Sec. 4 above).
@@
From a quantum chemistry perspective, if one considers the two BO electronic states of NO in one orientation as the degenerate "diabatic states" of interest, then the nuclear motion breaks this degeneracy, leading to the "adiabatic states" of interest in Fig. 5. However, the ee/$f$ states have identical energies at L = 0 because of the form of the splitting in Eq. 51, so that there are two different labeling schemes – Fig. 5a versus Fig. 5b. In a sense, one can argue that NO has a “conical intersection” 69, 2, 38 in momentum space, which provides a new interpretation of $\Lambda$-doubling (and a different view of the magnetic monopole derivative couplings discussed in Refs. 45, 10). Note that for the case of a diatomic molecule (as in Eq. 49), the conical intersection in phase space has co-dimension two (because only $L_{a}$ and $L_{b}$ enter the rotational Hamiltonian); more generally, in line with Ref. 19, we would expect the conical intersection around ${\bf L}=0$ to have co-dimension three. Note also that our analysis above of Eq. 51 is effectively quantitative insofar as a purely quantum calculation using $L_{a}=L_{b}=\frac{1}{2}(J+\frac{1}{2})$ will necessarily lead to an energy splitting proportional to a factor $2\cdot\frac{1}{2}(J+\frac{1}{2})\approx\sqrt{J(J+1)}$, which is in agreement with Eq. 51 above.
The $^{2}\Pi_{3/2}$ manifold
During the review process for this manuscript, a reviewer asked a very interesting question related to the scaling above. Namely, the reviewer asked us if a phase space method would recover the correct form for the $\Lambda$-splitting within the $^{2}\Pi_{3/2}$ manifold of NO. The question above is interesting for two reasons. First, the origin of the $\Lambda$-splitting for the $^{2}\Pi_{1/2}$ manifold is different than the origin of $\Lambda$-splitting for the $^{2}\Pi_{3/2}$ manifold. For the former, the dominant contribution is the double perturbation of $H_{\mathrm{SOC}}^{(1)}$ and the Coriolis potential $H^{(1)}_{\mathrm{Cor}}$. By contrast, for the latter, the lambda splitting has a different origin; here, the effect arises exclusively from the Coriolis potential, i.e. a second order perturbative treatment in $H^{(1)}_{\mathrm{Cor}}$ is required. As a result of this difference, the lambda splitting scales as $J$ for the former (see Fig. 4) and cubically for the latter. Obviously, one must wonder if a phase space approach can recover this difference? After all, if one were to follow Methods (i) and (ii) in Sec. 5, i.e. evaluating $\alpha$ at ${\bf P}=0$ and then diagonalize a matrix as in Sec. 4, one would presumably find a splitting that is always linear in $J$.
Second, a keen observer will note that, in order to recover the $\Lambda$-doubling results in Fig. 4 above, all that is required is that (i) one include SOC within a BO calculation and then (ii) one couple together the nuclear and electronic angular momentum through the $\hat{{\bf L}}^{e}$ operator, as in Eq. 42 above. In other words, one can in fact recover the $\Lambda$-doubling splitting of the $^{2}\Pi_{1/2}$ manifold without really ever invoking the phase space electronic structure Hamiltonian in Sec. 2; if one wishes to avoid exploring phase space potential energy surfaces, one can model $\Lambda$-doubling directly with the relativistic BO states by including $\mathbf{\hat{L}_{e}}$ alone (and without building the $\hat{\mathbf{\Gamma}}$ operator). That being said, just considering the scaling argument above, such an approach is clearly not appropriate for the $^{2}\Pi_{3/2}$ manifold. For the latter case, the only possibility is to invoke Method (iii), fully explore the phase space potential energy surfaces, and then investigate the energy gap at $L=n\hbar/2$. As a result, such an approach will inevitably require that we build $\hat{\Gamma}$ for ${\bf P}\neq 0$ (i.e. not just $\hat{{\bf L}}^{e}$) and whether or not PS can recover the correct $\Lambda$-splitting for the $^{2}\Pi_{3/2}$ manifold is in fact a very strenuous test of the method.
Figure 6: Energy splitting between the $e$ and $f$ states of the $^{2}\Pi_{3/2}$ manifold as a function of total angular momentum $J$. For this data, unlike Fig. 4, it is well known that the energy splitting requires a cubic (rather than linear) fit as a function of $J$. We compare phase space predictions, using a model parameter $\alpha$ derived from FCI with 1-electron SOC in the STO-3G basis with Method III (blue dots) versus experiment (black triangles). We further fit the “phase space” data to an analytic expression from standard perturbation theory, $\Delta E=k(J+\frac{1}{2})(J+\frac{3}{2})(J-\frac{1}{2})$, where $k$ is a fitted constant28 (dashed lines). Note that, as in Fig. 4 above, our phase space predictions are quite accurate, which is another strong confirmation of a PS approach.
As shown in Fig. 6, phase space theory does yield an accurate $\Lambda$-splitting as compared with experiment. Note that, in order to resolve the tiny absolute energy scale (on the order of $10^{-10}$ Hartree), we required a tight convergence threshold for the diagonalization. In Fig. 6, we also fit our data to an analytical energy expression from standard perturbation theory (valid under slow rotation and strong SOC) 28, namely $\Delta E\approx 7\times 10^{-6}(J+\frac{1}{2})(J+\frac{3}{2})(J-\frac{1}{2})$ cm-1 (see Fig. 6, blue dashed curve). Our results show a good fit, with a $R^{2}$ value of 0.9995. Altogether, the data in Fig. 6 again demonstrate the strong accuracy that can be found by solving electronic structure problems within a phase space framework.
@@ 7 Conclusions and Outlook
We have demonstrated that a simple phase space perspective on electronic structure theory provides a natural framework to capture the $\Lambda$-doubling in diatomic molecules. Thus, in line with previous discussions6, 11, we assert that the same beyond-BO physics that underlies one of the smallest microscopic diatomic phenomena is also responsible for the macroscopic Einstein de-Haas effect21, 44, 23 and, perhaps, chiral-induced spin selectivity9, 50, 25. Interestingly, the major result here has been the nonadiabatic splitting $\Delta E$ between different states of different angular momentum, where dynamically $\hbar/\Delta E$ corresponds to the tunneling time between degenerate states. In the context of the Einstein de Haas effect, one is less interested in parity eigenstates (e.g. ee/$f$), but rather in a basis with maximally aligned and opposite spin directions along the magnetic field direction. Nevertheless, the approach above should be able to isolate the tunneling time for a spin (or a collection of spins) to change orientation. In other words, whereas most treatments of Einstein de Haas phenomena use angular momentum conservation to predict the steady state nuclear rotational angular momentum that arises upon a total change in electronic spin angular momentum, the present approach based on phase space electronic structure theory will be able to go further and calculate transient effects: How long does it take for the spins to align? For the metal to reach a steady state of rotation? And of course, in the long run, how much energy is lost if we allow for frictional losses from nuclear vibrations?
@@
With regards to this last point, although in this work we have restricted our attention to electron coupling to purely rotational nuclear motion, the most interesting physics undoubtedly lies ahead as we consider larger molecules and materials, and will need to model both vibrations and rotations – and their coupled effect on electrons 14, 17, 16 – all on an equal footing. That being said, in the future, if we seek to model larger molecules or materials, it will be essential to efficiently build the $\hat{\mathbf{\Gamma}}$ operators – which will require some numerical tricks on our end and a more efficient implementation. Our results here were calculated in PySCF56, 57, and our group also has a preliminary code in the Q-Chem package22. Our strong feeling is that, given the success of the current manuscript in tandem with previous results in VCD20 and ROA59, we will soon be equipped to explore a very broad set of questions: what are the roles of chiral phonons in determining spin-dependent electron transfer rates 33? How should we best quantitatively describe spin relaxation41? Can we definitively prove the nature of the fundamental mechanisms behind different magnetic field effects53? Lastly, this manuscript has suggested that whenever the ground state has a degeneracy built on different electronic angular momenta, nuclear motion will lead to a small splitting. As such, might not the equations above (especially Eq. 49, which describes a gap opening due to off-diagonal electron-phonon coupling and is analogous to a Su–Schrieffer–Heeger (SSH)/Peierls Hamiltonian in the case of electron hopping55, 54) also apply to a metal and be a starting point for superconductivity calculations? When we move beyond a simple diatomic molecule, we can expect the number and the nature of interesting problems involving angular-momentum transfer to grow very rapidly.